# AdaptNC: Adaptive Nonconformity Scores for Uncertainty-Aware Autonomous Systems in Dynamic Environments

### v1

@@ 4 AdaptNC Method
Armed with our high-density region $\hat{R}_{t}$, we fit a convex hull shape template to cover these points. The parameters of this shape template are $\theta_{t}=\{A\in\mathbb{R}^{r\times p},b\in\mathbb{R}^{r}\}$, where $r$ is the number of facets of the hull. The non-conformity score is then defined based on the distance of a point $y$ from the convex hull defined by these shape templates:
$s(X_{t},y;\theta_{t})=\max_{j\in{1,\ldots,p}}A_{j}(y-h(X_{t}))-b_{j}$ (6)
We identify this convex hull of the points in $\hat{R}_{t}$ using the QuickHull algorithm (Barber et al., 1996).
@@
While we could directly use this score function in the threshold update step, as the underlying DtACI algorithm allows for changes in the score function, we pay an insiduous penalty for this. This penalty is expressed in the regret guarantee Theorem C.1, where the error is bounded by a term proportional to the square root of the difference between the true $\alpha^{*}$ in consecutive time steps. While this is normally small, changing the non-conformity score function parameters can result in a very large jump in the $\alpha^{*}$. Theoretical and empirical evidence for the necessity of the replay methodology is presented in Section 5.2 and Fig. 2.
In order to mitigate this difference, we replay the past $W$ data points with a freshly initialized set of expert weights and the new score function parameters $\theta_{t}$. This can be interpreted as a counterfactual evaluation of the conformal threshold update, where we identify the the value of $\alpha_{t}$ under the assumption that previous observations were scored under the parameters $\theta_{t}$ as opposed to other score parameters. Due to the exponential weight updates, the estimate of $\alpha_{t+1}$ is tied to the estimate of $\alpha_{t}$, which may have been made using a $\beta_{t}$ that is much different to the current one by virtue a change in the score parameter. The counterfactual Replay mitigates this problem by resetting the $\alpha_{t}$ update chain and evaluating past data through the lens of the new score parameters.
@@ 5 Theoretical Analysis
We will show that our method preserves the theoretical guarantees of DtACI, namely the regret and long-term coverage guarantees. We will build on this theory, and show that the score adaptation procedure we propose will also stabilize in the same setting as the long-term coverage guarantee. Finally, we will show a decomposition of distribution shift into the data-driven and score function driven parts, and provide evidence that this shift is observed in practice.
@@ 5.1 Coverage Guarantees
Our use of DtACI provides a regret guarantee for the adaptive miscoverage level selection procedure. The following remark is adapted from Gibbs & Candès (2024), and presented in full in Theorem C.1.
The Remark 5.2 states that the regret of our adaptive miscoverage level selection procedure is bounded by two terms: a term that decreases with increases in the size of the window $|W|$, and a term that depends on the total variation in the optimal miscoverage levels $\alpha^{*}_{t}$ over the window. This means that if the optimal miscoverage levels do not change too much over time, our procedure will perform nearly as well as if we had known the optimal levels in advance.
Assumption 5.1.
Let $\gamma_{max}\coloneqq\max_{1\leq i\leq k}\gamma_{t}$ and assume that $\gamma_{1}<\gamma_{2}<...<\gamma_{k}$ where $\gamma_{i+1}/\gamma_{i}\leq 2$ for all $1<i\leq k$, and that $\ell(\beta_{t},\theta)\coloneqq\alpha(\beta_{t}-\theta)-\min\{0,\beta_{t}-\theta\}$
@@
Then, for any window $W=[r,s]\in[T]$ and any sequence $\alpha^{*}_{r},\ldots,\alpha^{*}_{s}\in[0,1]$, and $\eta=\sqrt{\frac{\log(2k|W|)+1}{\sum_{t=r}^{s}\mathbb{E}[\ell(\beta_{t},\alpha_{t})^{2}]}}$ and $\sigma=1/(2|W|)$,
$\begin{split}\frac{1}{|W|}\sum_{t=r}^{s}\ell((\beta_{t},\bar{\alpha}_{t}))-\frac{1}{|W|}\sum_{t=r}^{s}\ell((\beta_{t},\alpha^{*}_{t}))\leq\\ O\left(\frac{\log(|W|)}{|W|}\right)+O\left(\sqrt{\frac{\sum_{t=r+1}^{s}\left|\alpha_{t}^{*}-\alpha_{t-1}^{*}\right|}{|W|}}\right).\end{split}$ (7)
The DtACI procedure has a randomized variant which provides the guarantee of long-term coverage (Gibbs & Candès, 2024, Theorem 6) approaching the target level Eq. 1. This theorem is presented in Theorem C.2, and it assumes that in a setting when the distribution stabilizes, we can send $\sigma,\eta\to 0$ over time. We show that under these same conditions, the non-conformity score function will also stabilize. The proof of the theorem is in Appendix Section C.3.
Theorem 5.3 (Score Function Stability).
Consider the same modified version of Algorithm 3 as in Theorem C.2. Let $\lim_{t\to\infty}\eta_{t}=\sigma_{t}=0$, and the number of data samples and Monte Carlo samples in Algorithm 4 $N,M\to\infty$. Then the non-conformity score functions will stabilize: $s(\cdot;\theta_{t})=s(\cdot;\theta_{t+1})$ as $t\to\infty$.
5.2 Replay
While changing the score function maintains the regret guarantee in Theorem C.1, the changes can have an outsize impact on the regret by virtue of producing changes in $\alpha^{*}$. The counterfactual replay mechanism proposed in Section 4 aims to mitigate this issue. In this section, we provide an intuitive explanation and empirical evidence in support.
Figure 2: This figure shows the change in the optimal threshold $\alpha_{t}^{*}$ based on scores from the initial and final distributions, $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$. The bottom pane shows the rate of distribution change, while the top shows the difference between optimal quantiles.
Proposition 5.4.
Let $\mathcal{D}_{t}$ and $\mathcal{D}_{t+1}$ be the joint distributions of inputs and observations $(X_{t},Y_{t})$ at times $t$ and $t+1$. Assume that the total variation distance ($\mathbf{TV}$) between the two distributions is bounded $\mathbf{TV}(\mathcal{D}_{t},\mathcal{D}_{t+1})\leq\delta$. For a fixed score function $s(\cdot)$, and a fixed score parameter $\theta$ let $S_{\theta,t}$ be the random variable $s(X,Y,\theta)$ where $(X,Y)\sim\mathcal{D}_{t}$, and $S_{\theta,t+1}$ be the random variable where $(X,Y)\sim\mathcal{D}_{t+1}$ Then, the $\mathbf{TV}$ distance between the score distributions will be less than $\delta$: $\mathbf{TV}(S_{\theta,t},S_{\theta,t+1})\leq\delta$
Proof.
This follows from the Data Processing Inequality (Polyanskiy & Wu, 2025, Theorem 7.4) with score function $s(\cdot;\theta)$ as channel, and $f$-divergence as $\mathbf{TV}$ distance. ∎
Remark 5.0.
The score distribution is a function of the underlying data and the score function used to process it. Assuming the shift in We can decompose the distribution shift as follows due to Proposition 5.4.
@@ 5.2 Replay
We are concerned mainly with the change that this distribution shift induces in the optimal $\alpha^{*}$. While we do not have theory for this, we provide experimental evidence in a controlled setting, where data is drawn from a mixture of two Gaussians $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, mixed by a changing weight $w_{t}$, we calculate the optimal $\alpha^{*}_{t}$ where the score function is the negative log-likelihood for each distribution. We can see in Fig. 2 that the distribution shift induces large changes in $\alpha^{*}$, which would induce large regret in the score update procedure. The theoretical maximum difference is 11, and the maximum observed difference was 0.7-0.7. Further details about this experiment are provided in Appendix D.1.
@@ 6 Experiments
Our objective is to demonstrate that jointly adapting the nonconformity score and its threshold enables tight uncertainty regions under distribution shift while preserving user-specified coverage. We evaluate this using global coverage, defined as $1-\frac{1}{T}\sum_{t=0}^{T}\operatorname{\textbf{err}}_{t}$, which measures recovery of the target coverage level, and the average coverage volume per covered timestep, which quantifies conservatism. To assess temporal behavior, we report a local coverage metric, defined by $1-\frac{1}{w}\sum_{t-\frac{w}{2}+1}^{t+\frac{w}{2}}\operatorname{\textbf{err}}_{t}$, a sliding-window average of size $w$ which is particularly informative under distribution shifts.
Figure 3: This figure illustrates the evolution of local coverage over a sliding window of 100 timesteps. For readability, an exponential moving average is shown, with the raw data in a lighter color. AdaptNC exhibits the lowest variability in local coverage among all methods, indicating stable recovery of tight uncertainty regions under distribution shift. In contrast, AdaptNC without Replay shows substantially higher variability, highlighting the role of the replay mechanism in mitigating coverage shocks induced by distribution shifts.
Baselines. We evaluate the proposed method against three baselines. (1) DtACI: DtACI adapts coverage over time by adjusting the prediction threshold but employs a fixed nonconformity score. Comparing against DtACI isolates the effect of score-function adaptation and enables the assessment of AdaptNC’s ability to recover tighter uncertainty regions under distribution shift. (2) Split Conformal Prediction: Standard split conformal prediction relies on the exchangeability assumption and does not account for distribution shifts. Comparison with this baseline highlights the importance of adaptivity in settings with distribution shift. (3) AdaptNC without Replay: As an ablation, we consider AdaptNC without Replay to examine the role of replay in stabilizing adaptation and mitigating coverage shock.
Figure 4: This figure illustrates the evolution of empirical coverage over the evaluation horizon relative to the target coverage level (shown in red). The proposed method consistently remains close to the desired coverage across time, whereas baseline methods exhibit pronounced periods of over- or under-coverage and, when recovery occurs, do so only after substantial deviation. This behavior highlights the flexibility of AdaptNC in maintaining calibrated coverage over time by adapting the nonconformity score function, rather than adjusting the threshold alone, enabling tighter uncertainty regions with sustained coverage.
6.1 Indoor Localization under Multipath Fading
In this case study, we evaluate the performance of AdaptNC in a setting with environmental shift. The case study examines an indoor localization scenario, where a robot continuously estimates its planar position using received signal strength indicator (RSSI) measurements from a small number of fixed WiFi routers. The robot operates in a bounded two-dimensional indoor environment with $4$ routers at known locations and follows double-integrator dynamics while executing random bounded motions. RSSI measurements are generated using a standard log-distance path-loss model with temporally correlated shadowing (Gudmundson, 1991) and small-scale fading (Jakes & Cox, 1994; Rappaport, 2010), capturing realistic wireless channel variability. Localization is performed using a fixed, model-based predictor that inverts RSSI measurements to distance estimates and combines iterative least-squares position estimation with an $\alpha$$\beta$ tracking filter. Importantly, the predictor neglects shadowing and fading effects and is tuned only under nominal conditions, with no learning or online adaptation at test time. As a result, a distribution shift emerges from the persistent temporal evolution of the wireless channel, which is not captured by the predictor’s underlying assumptions. Evaluation is conducted over a single rollout of 6000 time steps without resets, during which this mismatch accumulates over time, yielding a realistic and challenging benchmark for adaptive uncertainty quantification. Full simulation details are provided in Appendix D.2.
Figure 5: This figure visualizes the uncertainty regions produced by AdaptNC and DtACI at three representative timesteps ($t\in[720,1560,3120]$) in the multirotor tracking task, paired with the realized residual at each timestep. The results show that AdaptNC recovers markedly tighter uncertainty regions while maintaining coverage. In contrast, DtACI adapts the score threshold, constraining the region geometry and leading to misalignment with the residual distribution in the presence of structural distribution shifts.
Performance Evaluation and Comparative Analysis: As shown in Table 1, AdaptNC achieves a global coverage of $90.65\%$, which is closest to the target level of $90\%$ among all methods, while attaining the smallest volume per covered timestep. In contrast, DtACI and Split Conformal Prediction exhibit overcoverage with substantially larger uncertainty volumes, corresponding to $\mathbf{130\%}$ and $\mathbf{326\%}$ of AdaptNC’s volume, respectively, reflecting their limited ability to adapt to distribution shift. AdaptNC without replay similarly exhibits overcoverage and larger uncertainty regions, with volumes reaching $\mathbf{130\%}$ of AdaptNC’s. As further illustrated in Figures 3 and 4, the absence of replay leads to increased variability in local coverage, which propagates to oscillations in global coverage and necessitates more conservative uncertainty regions. These results highlight the importance of the replay mechanism, which stabilizes coverage under distribution shifts and enables the recovery of tighter, more accurate uncertainty regions.
6.2 Social Navigation with Policy Shift
In this case study, we consider a social navigation scenario in which multiple agents move through a shared environment while repeatedly navigating toward randomly assigned goals. Agent motion follows the social force model of Helbing & Molnár (1995), combining attractive forces toward goals with repulsive interactions from nearby agents and workspace boundaries, where interactions are limited to a finite collaboration radius. In our experiments, eight agents navigate a $10\,\mathrm{m}\times 10\,\mathrm{m}$ workspace over 6000 timesteps, receiving new goals upon reaching their current ones. A policy shift is introduced by gradually expanding the collaboration radius from $2\,\mathrm{m}$ to $5\,\mathrm{m}$, which alters how agents anticipate and respond to surrounding traffic over time. Trajectory prediction is performed using an LSTM trained under nominal conditions to predict the ego agent’s position five timesteps ahead from the past ten timesteps of observations, including the ego state, ego goal, and relative positions of other agents. At evaluation time, the expanding interaction radius induces a growing mismatch between the predictor’s training distribution and the evolving system behavior, as agents begin to slow down earlier or deviate to avoid more distant neighbors. This unmodeled policy shift gives rise to a distribution shift that degrades the coverage of non-adaptive conformal prediction methods, such as Split CP. Full details of the simulation setup are provided in Appendix D.3.
@@ 6.2 Social Navigation with Policy Shift
Method Global Cov. Avg. Volume $\downarrow$ Mean Local Cov. Global Cov. Avg. Volume $\downarrow$ Mean Local Cov. Global Cov. Avg. Volume $\downarrow$ Mean Local Cov.

### v2

@@ 4 AdaptNC Method
Given the estimated high-density region $\hat{R}_{t}$, AdaptNC fits a convex shape template to cover these points, in a manner similar to [25]. We parameterize the template by $\theta_{t}=\{A\in\mathbb{R}^{r\times p},b\in\mathbb{R}^{r}\}$, where $r$ is the number of facets. The resulting non-conformity score is: $s(X,y;\theta_{t})=\max_{j\in\{1,\ldots,r\}}\left(A_{j}(y-h(X))-b_{j}\right).$
Under this parameterization, the zero sublevel set of the score corresponds to the fitted convex template centered at the predictor output $h(X)$. We fit the convex hull of $\hat{R}_{t}$ using the QuickHull algorithm [2]. This produces convex prediction-region geometry, which is useful for downstream planning and control applications [19].
@@
Changing the score parameters alters the distribution of non-conformity scores. Consequently, the DtACI state prior to an update may remain calibrated to the previous score function. This mismatch induces coverage shock, a transient calibration error immediately following a score update. In the DtACI regret bound, this appears as a potentially large change in the hindsight-optimal miscoverage level $\alpha_{t}^{*}$ across consecutive timesteps (see Sections˜5.3 and E.1 for theoretical and empirical evidence ).
To mitigate coverage shock, AdaptNC replays the most recent $W$ observations after each score update. The replay procedure rescans these observations under the updated score parameters $\theta_{t}$, recomputes the corresponding quantiles $\beta_{t}$, and reruns the threshold updates. This can be interpreted as a counterfactual reconstruction of the DtACI state under the new score function. The resulting miscoverage level and threshold are then used for subsequent online predictions, effectively resetting the the threshold adaptation process under the new score distribution and reducing the abrupt calibration error induced by score updates.
@@ 5 Theoretical Analysis
AdaptNC modifies adaptive conformal prediction in two ways: it updates the nonconformity score online, and it replays recent observations after score updates to realign the ACI expert states with the updated score. These two mechanisms serve different goals. Score adaptation controls the geometry and efficiency of the prediction region, while the ACI expert updates control long-run coverage. Our theoretical analysis therefore separates these effects: we first study long-run stability of prediction-region size, then show that single-expert ACI coverage is preserved whenever replay induces only a vanishing average perturbation of the expert recursion. $\theta_{t}$.
5.1 Long-Run Score Function and Region Stability
We make a series of assumptions in the long-run stability and coverage analysis, which are analogous to the long-run regimes considered in prior work [11].
Assumption 5.1 (Weighted Residual Stability).
We assume that the weighted distribution of residuals $\mathcal{Z}_{t}$ converges in $f$-divergence to the steady state $\mathcal{Z}_{*}$ in the asymptote: $D_{\phi}(\mathcal{Z}_{t},\mathcal{Z}_{*})\to 0$ as $t\to\infty$. We additionally assume that the MCKDE parameters $M,N\to\infty$.
In AdaptNC, because the score function is fit from weighted residuals, the corresponding stability object is the weighted residual distribution used by the score-adaptation step. We require the regularity conditions for the HDR stability result in Assumption˜D.4. We also assume that the MCKDE procedure converges in Assumption˜C.1. The proof of Theorem˜5.2 can be found in Section˜D.2.
Theorem 5.2 (Score Function Stability).
[…]
$\sup_{(x,y)\in\mathcal{A}}|s_{K_{m}}(x,y)-s_{K_{\star}}(x,y)|\xrightarrow{p}0\quad\therefore\quad\sup_{(x,y)\in\mathcal{A}}|s_{K_{m+1}}(x,y)-s_{K_{m}}(x,y)|\xrightarrow{p}0.$
Corollary 5.3 (Stability of prediction-region size).
Suppose $\sup_{(x,y)\in\mathcal{A}}|s_{t}(x,y)-s_{\star}(x,y)|\leq\delta_{t}$ with $\delta_{t}\to 0$ in probability due to Theorem˜5.2. Under Assumption D.4, for every fixed threshold $q$,
$\left|\operatorname{Vol}(C_{t}(x;q))-\operatorname{Vol}(C_{\star}(x;q))\right|\leq L_{\mathrm{vol}}\delta_{t}\quad\textrm{in probability}.$
Hence the size of the prediction region stabilizes as the score function stabilizes in probability.
Thus, the score-stability result implies that the geometry and volume of the prediction regions stabilize around the limiting convex-template region. In this sense, the score-adaptation analysis is an efficiency result: once the residual geometry stabilizes, AdaptNC no longer incurs persistent changes in prediction region size due to score updates.
@@ 5.2 Coverage Guarantees
The long-term coverage guarantee of our method follows from a variation on the per-expert guarantee of ACI, which we present in Theorem˜5.4. This modification is necessary because the recursion used to prove the original ACI guarantee is perturbed by the score updates and replay mechanism. We show that the single-expert ACI guarantee is preserved whenever replay induces only a vanishing average perturbation of the expert recursion. We then give a sufficient condition for this perturbation to vanish: successive score updates must become small on replayed observations, and replayed observations must not concentrate near the conformal boundary.
Theorem 5.4 (Per-Expert Coverage Guarantee with AdaptNC).
Fix an expert $i\in\{1,\ldots,k\}$ with step size $\gamma_{i}>0$. At time $t$, let the expert’s prediction set be $\mathcal{C}_{t}^{i}=\mathcal{C}_{t}(X_{t};\theta_{t},q_{\mathcal{D}_{t}}(1-\alpha_{t}^{i})),$ and define its realized miscoverage indicator $\operatorname{\textbf{err}}_{t}^{i}=\mathds{1}\{Y_{t}\notin\mathcal{C}_{t}^{i}\}.$ Suppose the expert state evolves according to the perturbed ACI recursion $\alpha_{t+1}^{i}=\alpha_{t}^{i}+\gamma_{i}(\alpha-\operatorname{\textbf{err}}_{t}^{i})+\Delta_{t}^{i},$ where $\Delta_{t}^{i}$ is the perturbation caused by score updates and replay. Assume that the expert states remain uniformly bounded in the range $(0,1)$, then:
$\frac{1}{T}\sum_{t=1}^{T}|\Delta_{t}^{i}|\longrightarrow 0\implies\frac{1}{T}\sum_{t=1}^{T}\operatorname{\textbf{err}}_{t}^{i}\longrightarrow\alpha$ (6)
Theorem 5.5 (Averaged Replay Perturbation Bound).
Fix an ACI expert $i$ with step size $\gamma_{i}>0$. Let $\{t_{m}\}_{m\geq 1}$ be the score-update times, and let $\mathcal{R}_{m}$ be the replay window used at update $t_{m}$. Let $s_{m}^{-}$ and $s_{m}^{+}$ denote the score functions immediately before and after the update. Define $\varepsilon_{m}:=\sup_{r\in\mathcal{R}_{m}}\left|s_{m}^{+}(X_{r},Y_{r})-s_{m}^{-}(X_{r},Y_{r})\right|.$ Let $q_{m,r}^{i,-}$ and $q_{m,r}^{i,+}$ be the conformal thresholds used by expert $i$ at replay step $r$ under the pre-update and post-update scores. Suppose there exists $L_{q}<\infty$ such that $|q_{m,r}^{i,+}-q_{m,r}^{i,-}|\leq L_{q}\varepsilon_{m}$ for all $m$ and $r\in\mathcal{R}_{m}$. Suppose further that
$\frac{1}{T}\sum_{m:t_{m}\leq T}\sum_{r\in\mathcal{R}_{m}}\mathbf{1}\left\{\left|s_{m}^{-}(X_{r},Y_{r})-q_{m,r}^{i,-}\right|\leq(1+L_{q})\varepsilon_{m}\right\}\longrightarrow 0.$
Then the replay-induced perturbations satisfy $\frac{1}{T}\sum_{t=1}^{T}|\Delta_{t}^{i}|\longrightarrow 0.$
The second supposition requires that replayed observations do not accumulate in the shrinking band around the conformal boundary where a small score update can change the miscoverage indicator. Score stability makes the width of this band shrink, but it does not by itself rule out many replayed points lying near the boundary. The assumption can be interpreted as a regularity condition on the data distribution. Taken together, the preceding results establish the single-expert long-run coverage property underlying AdaptNC. Proofs are provided in Appendix˜D. Since the DtACI long-run coverage guarantee holds for arbitrary sequences, AdaptNC inherits this guarantee for any sequence of score functions.
5.3 Replay
We formally define coverage shock as the $\alpha^{*}$ shift term from Remark˜D.2: $\textrm{coverage shock}=\left|\alpha_{t}^{*}-\alpha_{t-1}^{*}\right|.$ Together with the term that varies inversely with the window size $|W|$, this quantity dominates the regret. Minimizing coverage shock, i.e., the change in the optimal miscoverage level across consecutive timesteps, will improve regret performance. However, changes to the NCSF can have an outsized impact on this difference, motivating the counterfactual replay mechanism. Here, we present an intuitive explanation paired with evidence from a toy experiment in support.
The hindsight-optimal $\alpha_{t}^{*}$ depends on the distribution of non-conformity scores at time $t$. We decompose changes in this distribution into those arising from new observations and those induced by updates to the NCSF. First, we note that replacing a single data point yields a bounded change in the score distribution in Proposition˜D.9. Next, we show that the distance caused by the change in NCSF parameter is not bounded by the data-distribution shift.
@@ 5.3 Replay
The score distribution depends on both the underlying data and the NCSF used to process it. We can decompose the distribution shift as follows due to Proposition˜D.9.
@@
5.4 Computational Complexity Analysis
The computational complexity is dominated by replay associated costs, proportional to the size of the replay window $W$.
Remark 5.7 (Computational Complexity of AdaptNC).
Let $W$ denote the window size, $k$ the number of experts, $M$ the number of Monte Carlo samples, $d$ the state dimension, $r$ the number of facets of the convex hull, and $n=|\hat{R}_{t}|$ the number of high-density points retained after thresholding. Under the assumption $d\leq 3$, the amortized per-timestep cost of AdaptNC is $O\!\left(W+k+MWd/t_{s}\right).$
We defer a detailed analysis of computational complexity to Section˜C.5, and a discussion in Section˜B.3
@@ 6 Experiments
We evaluate AdaptNC across three case studies capturing common forms of distribution shift in robotics and dynamical systems: (i) environmental shift from unmodeled sensor noise, (ii) policy-induced shift from changing interaction dynamics, and (iii) dynamics shift from gradual system degradation. Together, these settings provide a comprehensive testbed for adaptive uncertainty quantification under evolving conditions.
We present three experimental settings. In the Indoor Localization experiments, a robot performs localization using wireless signal strength measurements. Distribution shifts arise in this setting because the underlying localization model fails to account for realistic wireless channel dynamics which vary over time. In the Social Navigation setting, $8$ agents navigate a shared environment based on information from agents within a collaboration radius. A trajectory predictor seeks to estimate the future poses of the agents. The distribution shift in this scenario stems from a gradual increase in the collaboration radius. In the Multirotor Tracking experiment, a predictor seeks to estimate the future position of a multirotor tracking a reference trajectory while maintaining a fixed altitude, but encounters distribution shift due to gradual degradation in the multirotor actuators. These three experiments showcase AdaptNC’s benefits in settings due to environmental changes, policy shift in multi-agent settings, and actuator degradation. Further details are in Appendices E.2, E.3, and E.4.
For each case study, we evaluate all methods across $50$ different random seeds to assess performance variability and robustness. For each seed, evaluation is conducted over a single 6000-step rollout without resets, and results are aggregated across all seeds.
@@
Evaluation Metrics: (1) Global Coverage, defined as $1-\frac{1}{T}\sum_{t=0}^{T}\operatorname{\textbf{err}}_{t}$, which measures recovery of the target coverage level, (2) Average Coverage Volume per Covered Timestep, which quantifies conservatism. (3) Local Coverage metric, defined by $1-\frac{1}{w}\sum_{t-\frac{w}{2}+1}^{t+\frac{w}{2}}\operatorname{\textbf{err}}_{t}$, a sliding-window average of size $w$ that captures temporal behavior under distribution shift.
Baselines: We compare against four baselines. (1) DtACI: As the state-of-the-art method for online conformal prediction under distribution shift, it serves as the primary and most relevant baseline as it consistently outperforms prior approaches [11]. (2) Split Conformal Prediction: assumes exchangeability and does not account for distribution shift. (3) AdaptNC d without Replay: an ablation to assess the role of replay in stabilizing adaptation and mitigating coverage shock. (4) Shape-CP: [25] uses optimized but static nonconformity scores, highlighting the limitations of static calibration under distribution shift.
Table 1: This table reports the proposed evaluation metrics across all methods and case studies, averaged over 50 seeds. AdaptNC achieves coverage closest to the $90\%$ target while maintaining low uncertainty volumes. It also exhibits the most stable and robust behavior, as reflected by the low standard deviation of all metrics.